OpenStudy (anonymous):

Solve the quadratic equation 0 = x^2 + 6x + 9 . would it be 3,-3?

OpenStudy (kira_yamato):

x = -3

OpenStudy (anonymous):

how come just -3?

OpenStudy (kira_yamato):

x^2 + 6x + 9 = (x + 3)^2

OpenStudy (anonymous):

(-6 +/- sqrt(36-4*9)) / 2 36-4*9 = 0 also : x^2 +6x + 9 = x^2 + 3x + 3x + 9 = x(x+3) + 3(x+3) = (x+3)(x+3)

OpenStudy (anonymous):

ok